An electric irob of resistance 20 ohm draws a current of 5 amperes calculate the heat produced in 30 seconds
1
Expert's answer
2016-10-05T15:06:03-0400
For 30 seconds, stand out Q = I^2*R*t=5^2*20*30=15000 J